# AgenticSciML: Collaborative Multi-Agent Systems for Emergent Discovery in Scientific Machine Learning

### v1

@@ A.2.1 Problem Setup
A.2.2 Agents Setup
Table 2 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table 2: Agent configurations for Poisson L experiment
@@ A.2.3 Results
Figure 9: Evolution of solution tree for Poisson equation on L-shaped domain. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the combined PDE residual and boundary condition MSE. The best solution is colored in orange.
@@
Figure 10: Multi-agent evolution metrics for the Poisson equation on L-shaped domain. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
@@
Figure 11: Comparison of ground truth, root, and champion solutions for Poisson equation on L-shaped domain. (a) Ground truth solution from finite difference method. (b) Root solution, Score = $3.32\times 10^{-2}$. (c) Champion solution, Score = $3.58\times 10^{-5}$. The score shows the combined PDE residual and boundary condition MSE. The Champion solution exhibits significant improvement from the root.
A.2.4 Novelty of the Champion Solution
The agents strategized to decompose the solution into multiple components. The proposer agent’s initial plan is to decompose the target solution as $u(x,y)=u_{nn}(x,y)+u_{p}(x,y)+u_{s1}(x,y)+u_{s2}(x,y)$. Here $u_{nn}$ is a 5-layer MLP, $u_{p}=(x^{2}+y^{2})/4$ is an analytical particular solution satisfying $\nabla^{2}u_{p}=1$, and $u_{s1},u_{s2}$ are analytical singular basis functions $Ar^{2/3}\sin(\frac{2}{3}(\theta-\pi/2))$ and $Br^{4/3}\sin(\frac{4}{3}(\theta-\pi/2))$ with trainable amplitudes $A$ and $B$. To goal is to have the singular terms capture the singularity at the origin, while $u_{nn}$ learns the smooth remainder.
However, the engineer agent made a critical bug in the implementation: a PyTorch .detach() call cuts off gradient flow to $A$ and $B$, preventing them from being learned, so $A$ and $B$ are kept as the initial values $A=B=0$ in the final model. Therefore, this 4-way decomposition strategy is not fully realized in practice. The actual learned decomposition is $u(x,y)=u_{nn}(x,y)+u_{p}(x,y)$. The two components are visualized in Figure 12.
Figure 12: Decomposition of the champion solution showing the two active components. Left: $u_{p}=(x^{2}+y^{2})/4$, the particular solution satisfying $\nabla^{2}u_{p}=1$. Right: $u_{nn}$, the MLP-learned component.
@@ A.2.4 Novelty of the Champion Solution
The code implements n_uniform = int(4096 * 0.7) = 2867 uniform points and n_importance = 4096 - 2867 = 1229 importance-sampled points. During training, each epoch samples a fresh batch of collocation points according to this scheme. Figure 13 visualizes the two samples.
Figure 13: Uniform and importance sampling of collocation points. Left: Uniform sampling. Right: Importance sampling with power-law distribution $p(r)\propto r^{-0.7}$ concentrates points near the origin. During each epoch of training, 70% of collocation points are sampled uniformly and 30% from the importance sampler.
Finally, the agents improved the optimization by first training with Adam optimizer for 12,000 epochs using a CosineAnnealingLR scheduler. After Adam, it switches to L-BFGS training for 3,000 epochs, where the collocation points are fixded and do not change between epochs.
A.3 Solving Burger’s Equation using PINN
@@ A.3.1 Problem Setup
A.3.2 Agents Setup
Table 3 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table 3: Agent configurations for Burger’s PINN experiment
@@ A.3.3 Results
Figure 14: Comparison of ground truth, root, and champion solutions for Burger’s equation PINN. The score shows the combined PDE residual, initial condition, and boundary condition errors.
Figure 15: Evolution of solution tree for Burger’s equation PINN. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the combined PDE residual, initial condition, and boundary condition errors. The best solution is colored in orange.
@@
Figure 16: Multi-agent evolution metrics for Burger’s equation PINN. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
A.3.4 Novelty of the Champion Solution
The agents proposed an adaptive gradient-enhanced PINN based on the retrieved knowledge base entry “Gradient-Enhanced Physics-Informed Neural Networks (gPINN)” [yu2022gradient]. The network architecture is a 6-block ResNet with Fourier embedding. The input $(t,x)$ is first embedded as $t,\sin(2\pi x),\cos(2\pi x),x]$ to exploit the periodic boundary conditions.
@@ A.3.4 Novelty of the Champion Solution
Inspired by [mcclenny2023self], $\lambda(t,x)$ are per-point adaptive weights, which are parameterized by a small 3-layer MLP $\lambda_{\text{net}}$. For each point $(t,x)$, the network outputs $\lambda(t,x)=\text{clamp}(\text{softplus}(\lambda_{\text{net}}(t,x))\cdot\alpha,0.1,100)$ where $\alpha=1/\text{mean}(\lambda)$ normalizes the average weight to 1. The $\lambda_{\text{net}}$ is optimized via gradient ascent using a separate Adam optimizer that maximizes the weighted losses while the main model minimizes them.
@@
A.4 Antiderivative Operator Learning
@@ A.4.1 Problem Setup
A.4.2 Agents Setup
Table 4 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table 4: Agent configurations for antiderivative operator experiment
@@ A.4.3 Results
Figure 17: Evolution of solution tree for antiderivative operator learning. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the average relative L2 error over the validation set. The best solution is colored in orange.
@@
Figure 18: Multi-agent evolution metrics for antiderivative operator learning. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
Figure 19: Root solution predictions for antiderivative operator learning on the validation dataset where input functions are sampled using GRF with different length scales.
Figure 20: Champion solution predictions for antiderivative operator learning on the validation dataset where input functions are sampled using GRF with different length scales.
Figure 21: Error comparison for antiderivative operator learning. Left: Root solution error distribution. Right: Champion solution error distribution. Champion solution exhibits significant improvement from the root across all frequency categories.
A.4.4 Novelty of the Champion Solution
@@
Therefore, it decides to implement the unconventional choice of a linear branch network to enforce the linearity. Additionally, “physics-informed DeepONet” [wang2021learning] retrieved from the knowledge base inspires the incorporation of a physics-informed loss
@@
A.5 Multiple-input Operator Learning for Reaction-Diffusion Equation
This problem tests the agents’ ability to design a neural operator that takes multiple input functions. We follow the same setup as [jin2022mionet], but this work is not included in the knowledge base.
A.5.1 Problem Setup
@@
A.5.2 Agents Setup
Table 5 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table 5: Agent configurations for multi-input operator experiment
@@ A.5.3 Results
Figure 22: Evolution of solution tree for multi-input operator learning for reaction-diffusion equation. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the average relative L2 error over the validation set. The best solution is colored in orange.
@@
Figure 23: Multi-agent evolution metrics for multi-input operator learning for reaction-diffusion equation. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
Figure 24: Root solution predictions for multi-input operator learning on the validation dataset. The contours and magnitudes of the prediction are not accurate compared to the ground truth.
Figure 25: Champion solution predictions for multi-input operator learning on the validation dataset. The prediction is very close to the ground truth.
A.5.4 Novelty of the Champion Solution
@@
Inspired by a retrieved knowledge base entry on the training strategy of U-FNO for capturing sharp saturation fronts in CO2 storage simulations [wen2022u], the proposer introduced a derivative loss term to encourage the model to learn solutions more accurately in regions of high spatial variation. The spatial derivative $\partial u/\partial x$ is approximated using a central finite difference scheme $(\partial u/\partial x)_{i}\approx(u_{i+1}-u_{i-1})/(2\Delta x)$. The derivative loss is computed as the relative loss
@@
A.6 Reconstruction of 2D Cylinder Wake Vorticity Field from Sparse Observations
This problem evaluates the agents’ capability to design a model architecture that can reconstruct a physical field from sparse and noisy sensor measurements. We use the set up and dataset from [zhao2024recfno] and [fukami2021global], but these works are not included in the agents’ knowledge base.
A.6.1 Problem Setup
@@
A.6.2 Agents Setup
Table 6 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table 6: Agent configurations for cylinder reconstruction experiment
@@ A.6.3 Results
Figure 26: Evolution of solution tree for cylinder wake vorticity field reconstruction from sparse noisy observations. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the relative L2 error on the validation set. The best solution is colored in orange.
@@
Figure 27: Multi-agent evolution metrics for cylinder wake vorticity field reconstruction from sparse noisy observations. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
@@
Figure 28: Comparison of ground truth, root, and champion solutions for cylinder wake vorticity field reconstruction.
A.6.4 Novelty of the Champion Solution
The solution combines the U-FNO architecture [wen2022u] with a bandlimit-preserving filter in the U-Net component inspired by Convolutional Neural Operators (CNO) [raonic2023convolutional] to mitigate aliasing effects during decoder upsampling. This strategy is different from the ones in [zhao2024recfno] and [fukami2021global] which inspired this benchmark problem.
@@ A.6.4 Novelty of the Champion Solution
Appendix B Knowledge Base
B.1 Construction of the Knowledge Base
@@
An example knowledge base entry is provided in Figure 29.
Figure 29: Example knowledge base entry and the JSON index summarizing method descriptions and file paths.
B.2 List of Knowledge Base Entries
@@
Table 7: Knowledge Base Entries: 70 Scientific Machine Learning Methods

### v2

@@ 4 Summary and Discussion
S1.2.2 Agents Setup
Table S3 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table S3: Agent configurations for Poisson L experiment
@@
Figure S4: Evolution of solution tree for Poisson equation on L-shaped domain. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the combined PDE residual and boundary condition MSE. The best solution is colored in orange.
@@
Figure S5: Multi-agent evolution metrics for the Poisson equation on L-shaped domain. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
@@
Figure S6: Comparison of ground truth, root, and champion solutions for Poisson equation on L-shaped domain. (a) Ground truth solution from finite difference method. (b) Root solution, Score = $3.32\times 10^{-2}$. (c) Champion solution, Score = $3.58\times 10^{-5}$. The score shows the combined PDE residual and boundary condition MSE. The Champion solution exhibits significant improvement from the root.
S1.2.4 Novelty of the Champion Solution
The agents planned to decompose the solution into multiple components. The proposer agent’s initial plan is to decompose the target solution as $u(x,y)=u_{nn}(x,y)+u_{p}(x,y)+u_{s1}(x,y)+u_{s2}(x,y)$. Here $u_{nn}$ is a 5-layer MLP, $u_{p}=(x^{2}+y^{2})/4$ is an analytical particular solution satisfying $\nabla^{2}u_{p}=1$, and $u_{s1},u_{s2}$ are analytical singular basis functions $Ar^{2/3}\sin(\frac{2}{3}(\theta-\pi/2))$ and $Br^{4/3}\sin(\frac{4}{3}(\theta-\pi/2))$ with trainable amplitudes $A$ and $B$. The goal is to have the singular terms capture the singularity at the origin, while $u_{nn}$ learns the smooth remainder.
However, the engineer agent made a critical bug in the implementation: a PyTorch .detach() call cuts off gradient flow to $A$ and $B$, preventing them from being learned, so $A$ and $B$ are kept as the initial values $A=B=0$ in the final model. Therefore, this 4-way decomposition strategy is not fully realized in practice. The actual learned decomposition is $u(x,y)=u_{nn}(x,y)+u_{p}(x,y)$. The two components are visualized in Figure S7.
Figure S7: Decomposition of the champion solution showing the two active components. Left: $u_{p}=(x^{2}+y^{2})/4$, the particular solution satisfying $\nabla^{2}u_{p}=1$. Right: $u_{nn}$, the MLP-learned component.
@@
The code implements n_uniform = int(4096 * 0.7) = 2867 uniform points and n_importance = 4096 - 2867 = 1229 importance-sampled points. During training, each epoch samples a fresh batch of collocation points according to this scheme. Figure S8 visualizes the two samples.
Figure S8: Uniform and importance sampling of collocation points. Left: Uniform sampling. Right: Importance sampling with power-law distribution $p(r)\propto r^{-0.7}$ concentrates points near the origin. During each epoch of training, 70% of collocation points are sampled uniformly and 30% from the importance sampler.
Finally, the agents improved the optimization by first training with Adam optimizer for 12,000 epochs using a CosineAnnealingLR scheduler. After Adam, it switches to L-BFGS training for 3,000 epochs, where the collocation points are fixed and do not change between epochs.
S1.2.5 Token and Cost Analysis
Table S4: Token and cost analysis for Poisson Equation on L-shaped Domain (PINN, 6 iterations). GPU training time dominates at 5.6 h vs. 1.7 h LLM time; the proposer accounts for 44% of LLM cost.
Agent (Model) API Calls Tokens (M) Cost (\$) Time (s)
Overall
   Total 156 5.54 10.68 26,413
   LLM (multi-agent) 156 5.54 10.68 6,174
   GPU Training 20,239
Agent Breakdown
   Proposer (Gemini-2.5-Pro) 57 1.71 (31%) 4.72 (44%) 2,723 (44%)
   Analyst (Gemini-2.5-Pro) 19 2.12 (38%) 3.31 (31%) 757 (12%)
   Engineer (Claude-Haiku) 21 0.47 (8%) 1.20 (11%) 1,022 (17%)
   Retriever (Gemini-2.5-Pro) 19 0.50 (9%) 1.08 (10%) 546 (9%)
   Critic (GPT-4o-Mini) 38 0.72 (13%) 0.36 (3%) 1,091 (18%)
   Debugger (GPT-4o-Mini) 2 0.03 (0%) 0.01 (0%) 35 (1%)
[…]
S1.3 Solving Burger’s Equation using PINN
@@
S1.3.2 Agents Setup
Table S5 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table S5: Agent configurations for Burger’s PINN experiment
@@
Figure S9: Comparison of ground truth, root, and champion solutions for Burger’s equation PINN. The score shows the combined PDE residual, initial condition, and boundary condition errors.
Figure S10: Evolution of solution tree for Burger’s equation PINN. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the combined PDE residual, initial condition, and boundary condition errors. The best solution is colored in orange.
@@
Figure S11: Multi-agent evolution metrics for Burger’s equation PINN. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
S1.3.4 Novelty of the Champion Solution
The agents proposed an adaptive gradient-enhanced PINN based on the retrieved knowledge base entry “Gradient-Enhanced Physics-Informed Neural Networks (gPINN)” [51]. The network architecture is a 6-block ResNet with Fourier embedding. The input $(t,x)$ is first embedded as $t,\sin(2\pi x),\cos(2\pi x),x]$ to exploit the periodic boundary conditions.
@@
Inspired by [31], $\lambda(t,x)$ are per-point adaptive weights, which are parameterized by a small 3-layer MLP $\lambda_{\text{net}}$. For each point $(t,x)$, the network outputs $\lambda(t,x)=\text{clamp}(\text{softplus}(\lambda_{\text{net}}(t,x))\cdot\alpha,0.1,100)$ where $\alpha=1/\text{mean}(\lambda)$ normalizes the average weight to 1. The $\lambda_{\text{net}}$ is optimized via gradient ascent using a separate Adam optimizer that maximizes the weighted losses while the main model minimizes them.
@@
S1.3.5 Token and Cost Analysis
Table S6: Token and cost analysis for Burgers’ Equation PINN (physics-informed neural network, 3 iterations). The Grok-4 proposer generates 31% of tokens at only 6% of cost; the engineer accounts for 40% of LLM cost.
Agent (Model) API Calls Tokens (M) Cost (\$) Time (s)
Overall
   Total 79 1.60 2.07 5,198
   LLM (multi-agent) 79 1.60 2.07 2,101
   GPU Training 3,097
Agent Breakdown
   Proposer (Grok-4) 24 0.50 (31%) 0.12 (6%) 481 (23%)
   Analyst (Gemini-2.5-Pro) 7 0.17 (11%) 0.45 (22%) 275 (13%)
   Engineer (Claude-Haiku) 15 0.33 (21%) 0.83 (40%) 647 (31%)
   Retriever (Gemini-2.5-Pro) 9 0.22 (14%) 0.49 (23%) 253 (12%)
   Critic (GPT-4o-Mini) 16 0.30 (19%) 0.15 (7%) 347 (17%)
   Debugger (GPT-4o-Mini) 8 0.08 (5%) 0.04 (2%) 97 (5%)
The Grok-4 proposer generates a large token volume (31%) at very low per-token cost (6% of total), while the engineer accounts for 40% of total LLM cost. The debugger is called 8 times across only 3 iterations, indicating that the generated code required notable repair. At \$2.07 total LLM cost this is the least expensive run.
S1.4 Antiderivative Operator Learning
@@
S1.4.2 Agents Setup
Table S7 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table S7: Agent configurations for antiderivative operator experiment
@@
Figure S12: Evolution of solution tree for antiderivative operator learning. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the average relative L2 error over the validation set. The best solution is colored in orange.
@@
Figure S13: Multi-agent evolution metrics for antiderivative operator learning. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
Figure S14: Root solution predictions for antiderivative operator learning on the validation dataset where input functions are sampled using GRF with different length scales.
Figure S15: Champion solution predictions for antiderivative operator learning on the validation dataset where input functions are sampled using GRF with different length scales.
Figure S16: Error comparison for antiderivative operator learning. Left: Root solution error distribution. Right: Champion solution error distribution. Champion solution exhibits significant improvement from the root across all frequency categories.
S1.4.4 Novelty of the Champion Solution
@@
Therefore, it decides to implement the unconventional choice of a linear branch network to enforce the linearity. Additionally, “physics-informed DeepONet” [46] retrieved from the knowledge base inspires the incorporation of a physics-informed loss
@@
S1.4.5 Token and Cost Analysis
[…]
Agent (Model) API Calls Tokens (M) Cost (\$) Time (s)
Overall
   Total 130 3.29 6.75 22,906
   LLM (multi-agent) 130 3.29 6.75 4,361
   GPU Training 18,545
Agent Breakdown
   Proposer (Gemini-2.5-Pro) 45 1.24 (38%) 3.36 (50%) 1,939 (44%)
   Analyst (Gemini-2.5-Pro) 15 0.70 (21%) 1.39 (21%) 597 (14%)
   Engineer (Claude-Haiku) 20 0.38 (12%) 0.91 (13%) 747 (17%)
   Retriever (Gemini-2.5-Pro) 15 0.37 (11%) 0.80 (12%) 391 (9%)
   Critic (GPT-4o-Mini) 30 0.54 (16%) 0.26 (4%) 592 (14%)
   Debugger (GPT-4o-Mini) 5 0.05 (2%) 0.03 (0%) 94 (2%)
[…]
S1.5 Multiple-input Operator Learning for Reaction-Diffusion Equation
This problem tests the agents’ ability to design a neural operator that takes multiple input functions. We follow the same setup as [18], but this work is not included in the knowledge base.
S1.5.1 Problem Setup
@@
S1.5.2 Agents Setup
Table S9 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table S9: Agent configurations for multi-input operator experiment
@@
Figure S17: Evolution of solution tree for multi-input operator learning for reaction-diffusion equation. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the average relative L2 error over the validation set. The best solution is colored in orange.
@@
Figure S18: Multi-agent evolution metrics for multi-input operator learning for reaction-diffusion equation. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
Figure S19: Root solution predictions for multi-input operator learning on the validation dataset. The contours and magnitudes of the prediction are not accurate compared to the ground truth.
Figure S20: Champion solution predictions for multi-input operator learning on the validation dataset. The prediction is very close to the ground truth.
S1.5.4 Novelty of the Champion Solution
@@
Inspired by a retrieved knowledge base entry on the training strategy of U-FNO for capturing sharp saturation fronts in CO2 storage simulations [48], the proposer introduced a derivative loss term to encourage the model to learn solutions more accurately in regions of high spatial variation. The spatial derivative $\partial u/\partial x$ is approximated using a central finite difference scheme $(\partial u/\partial x)_{i}\approx(u_{i+1}-u_{i-1})/(2\Delta x)$. The derivative loss is computed as the relative loss
@@
S1.5.5 Token and Cost Analysis
Table S10: Token and cost analysis for Multi-input Operator Learning for Diffusion-Reaction PDE (neural operator, 8 iterations). GPU training time dominates at 10.7 h vs. 2.1 h LLM time; the Gemini-2.5-Flash model enables cost-effective scaling to 8 iterations at \$6.73 total.
Agent (Model) API Calls Tokens (M) Cost (\$) Time (s)
Overall
   Total 204 8.10 6.73 46,248
   LLM (multi-agent) 204 8.10 6.73 7,713
   GPU Training 38,535
Agent Breakdown
   Proposer (Gemini-2.5-Flash) 72 2.68 (33%) 2.72 (40%) 3,644 (47%)
   Analyst (Gemini-2.5-Flash) 18 2.96 (36%) 1.10 (16%) 502 (7%)
   Engineer (Claude-Haiku) 33 0.80 (10%) 1.96 (29%) 1,597 (21%)
   Retriever (Gemini-2.5-Flash) 24 0.70 (9%) 0.49 (7%) 573 (7%)
   Critic (GPT-4o-Mini) 48 0.85 (11%) 0.42 (6%) 1,277 (17%)
   Debugger (GPT-4o-Mini) 9 0.11 (1%) 0.05 (1%) 122 (2%)
At 8 iterations and 204 LLM calls, this is the longest run; GPU training time (10.71 h) far exceeds LLM API time (2.14 h). The analyst incurs 36% of all tokens — the highest analyst share — due to multimodal outputs for the multi-input setting. Despite the longest run length, total LLM cost (\$6.73) remains moderate, enabled by the lightweight Gemini-2.5-Flash model.
S1.6 Reconstruction of 2D Cylinder Wake Vorticity Field from Sparse Observations
This problem evaluates the agents’ capability to design a model architecture that can reconstruct a physical field from sparse and noisy sensor measurements. We use the set up and dataset from [54] and [9], but these works are not included in the agents’ knowledge base.
S1.6.1 Problem Setup
@@
S1.6.2 Agents Setup
Table S11 summarizes the agent configurations and evolutionary parameters used in this experiment.
Table S11: Agent configurations for cylinder reconstruction experiment
@@
Figure S21: Evolution of solution tree for cylinder wake vorticity field reconstruction from sparse noisy observations. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the relative L2 error on the validation set. The best solution is colored in orange.
@@
Figure S22: Multi-agent evolution metrics for cylinder wake vorticity field reconstruction from sparse noisy observations. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
@@
Figure S23: Comparison of ground truth, root, and champion solutions for cylinder wake vorticity field reconstruction.
S1.6.4 Novelty of the Champion Solution
The solution combines the U-FNO architecture [48] with a bandlimit-preserving filter in the U-Net component inspired by Convolutional Neural Operators (CNO) [38] to mitigate aliasing effects during decoder upsampling. This strategy is different from the ones in [54] and [9] which inspired this benchmark problem.
@@
S1.6.5 Token and Cost Analysis
Table S12: Token and cost analysis for Cylinder Wake Flow Reconstruction (inverse problem, 5 iterations). The engineer accounts for 71% of LLM cost due to the complexity of the reconstruction pipeline; the debugger is invoked 31 times — the most of any experiment.
Agent (Model) API Calls Tokens (M) Cost (\$) Time (s)
Overall
   Total 182 4.99 4.88 27,929
   LLM (multi-agent) 182 4.99 4.88 5,390
   GPU Training 22,539
Agent Breakdown
   Proposer (Grok-4) 45 1.14 (23%) 0.27 (5%) 1,008 (19%)
   Analyst (Gemini-2.5-Flash) 15 0.90 (18%) 0.41 (8%) 305 (6%)
   Engineer (Claude-Haiku) 46 1.40 (28%) 3.44 (71%) 2,769 (51%)
   Retriever (Gemini-2.5-Flash) 15 0.46 (9%) 0.30 (6%) 345 (6%)
   Critic (GPT-4o-Mini) 30 0.67 (14%) 0.30 (6%) 647 (12%)
   Debugger (GPT-4o-Mini) 31 0.42 (8%) 0.16 (3%) 316 (6%)
The engineer accounts for 71% of total LLM cost and 51% of LLM time across 46 calls — a sharp reversal of the proposer-dominant pattern in all other experiments — reflecting the difficulty of setting up the inverse reconstruction pipeline. The debugger is invoked 31 times, confirming the cylinder codebase required the most iterative repair. The Grok-4 proposer again generates a large token share (23%) at minimal cost (5%), consistent with the Burgers’ experiment.
S2 Knowledge Base
S2.1 Construction of the Knowledge Base
@@
An example knowledge base entry is provided in Figure S24.
Figure S24: Example knowledge base entry and the JSON index summarizing method descriptions and file paths.
S2.2 Ablation Studies of the Knowledge Base
We evaluate KB sensitivity on one controlled benchmark: discontinuous function approximation (Supplementary Section S1.1). The objective is to test whether performance depends on retrieving relevant KB entries, and how the multi-agent reasoning trajectory changes when KB support is removed or made irrelevant.
Methodology. We ran three settings with all non-KB factors fixed:
  • Full KB: standard retrieval from the 70-entry KB.
  • No KB: retriever disabled (no KB context provided to proposer/critic).
  • Random KB: retriever returns a random KB entry each mutation.
All three runs use the same root solution (solution_0), the same evaluation contract (evaluate.py, guidelines.md), the same agent configuration in Table S1, and the same evolutionary parameters (6 iterations, mutation batch size 4). For this benchmark, training is done on train_data.npz; the score is always computed by the fixed evaluator on held-out val_data.npz. All three settings start from the same root MSE, 0.2828776240.
Table S13: KB ablation results for discontinuous function approximation (lower is better).
Setting Best solution Best MSE
Full KB solution_0100 0.0014631104
No KB solution_00101 0.0034207932
Random KB solution_00002 0.0303153563
Relative to Full KB, No-KB is $2.338\times$ higher error and Random-KB is $20.720\times$ higher error.
Table S14: Round-3 proposer trajectory comparison across KB settings (function approximation only).
Setting Best lineage Observed method trajectory
Full KB 0 $\rightarrow$ 01 $\rightarrow$ 010 $\rightarrow$ 0100 Trajectory converges to a stable MoE strategy: feature-augmented adaptive activations (01) $\rightarrow$ explicit two-expert gating (010) $\rightarrow$ bounded trainable gate sharpness and interface-focused weighting (0100). Retrieved entries are directly relevant, and later rounds refine rather than replace the core mechanism.
No KB 0 $\rightarrow$ $0$ $\rightarrow$ 001 $\rightarrow$ 0010 $\rightarrow$ 00101 Agents still discover domain-decomposition and gating ideas from internal reasoning, but with more trial-and-error pivots: soft temperature gate $\rightarrow$ hard detached gate $\rightarrow$ Gumbel-softmax routing $\rightarrow$ rollback to simpler asymmetric experts with a stable sigmoid gate and explicit jump term. Improvement is substantial, but less efficient and less stable than Full KB.
Random KB 0 $\rightarrow$ $0$ $\rightarrow$ 000 $\rightarrow$ $00$ $\rightarrow$ 00002 Retrieved papers are frequently off-task (e.g., PINN/DeepONet/GNOT/ev-NSFnet entries from different PDE/operator regimes). The trajectory is less coherent: continuity-heavy and auxiliary-complex detours are introduced, then partially removed. Final recovery comes from simplifying back to a cleaner MoE-like formulation, but quality remains much worse than Full KB.
These results indicate that success is not purely dependent on retrieving one exact “right” paper: even without KB, the system still iteratively improves candidate solutions from $0.2829$ to $0.00342$ through multi-agent reasoning alone, although the final result is less optimal than Full KB. At the same time, performance is strongly correlated with KB relevance. Relevant retrieval helps the proposal-critic loop converge to a coherent method and best score, while random/irrelevant retrieval degrades outcome quality and can misguide agent reasoning toward less relevant design choices.
S2.3 List of Knowledge Base Entries
@@
Table S15: Knowledge Base Entries: 70 Scientific Machine Learning Methods